# LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data

### v1

@@ 3.4.1. Algorithms
LOTUS’ implementation builds on and generalizes the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n. d.]; lan, [n. d.]). These implementations primarily leverage one of two aggregation patterns: either a fold pattern, which produces a sequential, linear pass over the data while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains.
3.4.2. Discussion
By default, LOTUS’ aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024). However, LOTUS’ partition function can be used to override the default functionality and generalizes the information flow of prior implementation by allowing users to specify arbitrary groupings over the data. Each unique group specified by the partition function will be aggregated with as few LM invocations as possible according to the model context length, before being merged with other groups. This allows the user to perform a fold pattern, as well as arbitrary user-defined patterns.
3.4.3. Optimizations
Preliminary results show that achieving high-quality semantic-aggregations is non-trivial and sensitive to the document’s input ordering and grouping. We see qualitative evidence of this in a summarization task of 50 ArXiv paper abstracts, which we show in Figure 21. We contrast the summarization results of a naive semantic aggregation (Figure 22) performed over the paper abstracts with the semantic aggregation performed using a partitioning function based on semantic similarity (Figure 23). The semantic aggregation using the partitioner first clusters documents based on semantic similarity, then aggregates each cluster before performing aggregations across clusters. As the figures show the two methods result in significantly different summaries. While the latter demonstrates qualitatively more cohesion and effectively abstracts general themes across papers, the former omits details and tends to list low-level details from individual papers rather than capturing higher-level themes. We leave a quantitative study of this to future work and believe that semantic aggregations create a rich design space for optimization.
3.5. sem_map & sem_extract
Both LOTUS’ semantic map and extract run batched LM calls over a set of rows, prompting the LM with the user’s arbitrary natural-language expression to generate a new column. Both functions can be fully parallelized over rows, and LOTUS implements efficient batched inference with vLLM (Kwon et al., 2023). To implement sem_extract, LOTUS prompts the model to answer the user’s langex with direct quotes, and the system then verifies that these snippets returned by the LM match the reference text.
3.6. Semantic Indexing
LOTUS supports several different algorithms for its semantic index.
To generate the semantic similarity index, LOTUS’ sim_index operator first batch processes the user-specified column to generate semantic embeddings with the configured retriever model, then constructs a vector index over the semantic embeddings. The current implementation uses FAISS’ flat index by default and writes the index locally to disk. User’s can specify the local file path to store the index, and optionally specify alternative vector indices, such as hierarchical navigable small worlds (HNSW) (Malkov and Yashunin, 2018), inverted indices (IVF) (Baranchuk et al., 2018; Ge et al., 2014; Jégou et al., 2011; Johnson et al., 2017), and locality sensitive hash indices (LSH) (Andoni and Indyk, 2008; Andoni and Razenshteyn, 2015; Indyk and Motwani, 1998; Jafari et al., 2020; Li et al., 2020; Liu et al., 2021; Lu and Kudo, 2020; Park et al., 2015; Sundaram et al., 2013; Zheng et al., 2020; Andoni et al., 2015; Gionis et al., 1999; Gong et al., 2020).
In the future, we envision additionally supporting a wider variety of embeddings stores and indices (Patel et al., 2024; Wang et al., 2021; vec, [n. d.]; noa, 2023b; chr, [n. d.]; mos, [n. d.]).
LOTUS’ semantic search, similarity join and cluster operations all leverage the similarity index in their implementation. sem_search first embeds the user’s query string using the configured retriever model, then performs an efficient top-k search using the loaded FAISS index over the search column. sem_sim_join similarly performs a top-k search using the loaded FAISS index over the right key. Here the right-key may be a column of natural-language text, which LOTUS will embed using the retriever model on-the-fly, or an indexed column, for which LOTUS will load the previously-generated embeddings. The right-key embeddings then serve as the queries to perform batched search over the search column, given by the left key. Lastly, LOTUS implements sem_cluster_by using FAISS optimized kmeans library to cluster the user-specified column. Here, the column must be previously indexed, and LOTUS uses the generated embeddings to perform the clustering with the user-specified number of cluster centroids.
@@ 4. Evaluation
We now evaluate LOTUS’ programmability and efficiency through three diverse applications: fact-checking (Section 4.1), extreme multi-label classification (Section 4.2), and search and ranking (Section 4.3). For each of these applications, we see that state-of-the-art quality results are achievable with low development overhead, using LOTUS programs with a few lines of code and a few or even one semantic operator. In addition, our results demonstrate interesting implementation and optimization choices introduced by semantic operators. Specifically, we find that:
On the FEVER dataset (Thorne et al., 2018) for fact-checking, LOTUS programs can reproduce FacTool (Chern et al., 2023), a recent state-of-the-art pipeline, in few lines of code, and implement a new pipeline with a simple change of operators that improves accuracy by , while offering lower execution time.
On the BioDEX dataset (D’Oosterlinck et al., 2023) for the extreme multi-label classification task, LOTUS reproduces state-of-the art result quality (D’Oosterlinck et al., 2024) with it’s join operator, while providing an efficient algorithm that runs up to faster than a naive join, demonstrating the power of LOTUS’ declarative interface.
On the SciFact dataset (Thakur et al., 2021) and two newly-constructed paper datasets for the search and ranking application, LOTUS’s semantic top- operator achieves higher nDCG@10 than a vanilla retriever and re-ranker, while also running faster than alternative LM-based ranking methods used by prior works (Qin et al., 2024).
Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama 3 models (int, [n. d.]), with a batch size of 64 running on vLLM (Kwon et al., 2023). For our experiments that use OpenAI’s GPT models (ope, [n. d.]), we run with 64-way thread parallelism.
@@ 4.1. Application: Fact-Checking
4.1.1. Dataset
We evaluate on FEVER (Thorne et al., 2018), a claim verification dataset. We use the development dataset, which contains about 38,000 total claims, of which we sample 500 for our evaluation. Each claim is labeled with one of three labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”, and the task is to correctly determine the label of each claim, leveraging evidence from a corpus of million Wikipedia articles. We merge the latter two labels in to a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation.
4.1.2. Baselines
FacTool (Chern et al., 2023) is a recent research work that proposes a multi-step pipeline for fact-checking involving, claim extraction, query generation, tool querying, evidence collection, and verification. We use FactTool’s open source codebase (gai, [n. d.]) to measure its performance. FactTool’s pipeline, by default, performs retrieval with a Google Search API (cus, [n. d.]). We evaluate the pipeline with both the default retrieval API, and alternatively test with a ColBERT (Khattab and Zaharia, 2020) index over the document corpus to perform retrieval. We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the implemented LOTUS programs.
| Method | Accuracy | Execution Time, batched (s) | Execution Time, no batching (s) | LoC |
| FacTool | 83.5 | N/A | 1,174.8 | ¿ 750 |
| LOTUS-Factool | 90.0 | 111.48 | 979.50 | ¡ 50 |
| LOTUS-fact-filter | 90.5 | 64.28 | 206.45 | ¡ 50 |
| LOTUS-fact-filter (+ cascades) | 93 | 34.09 | 168.24 | ¡ 50 |
| LOTUS-fact-join | 86.5 | 2,394.22 | 12,923.08 | ¡ 50 |
4.1.3. LOTUS Programs
We compose several intuitive LOTUS programs, each in less than 50 lines of code. For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama 70B as the primary LM, and Llama 8B and TinyLlama for cascade optimizations.
@@ 4.1.3. LOTUS Programs
The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the query pipeline, with LOTUS’ semantic filter operation. We use 3 demonstrations for the filter. Figure 25 shows the pseudocode for this program. We evaluate this program with and without model cascades for the semantic filter operation.
@@ 4.1.4. Results
Table LABEL:tab:factchecking demonstrates the powerful abstraction that LOTUS provides, allowing programmers to quickly write and test programs that compose simple operators to obtain state-of-the-art results. We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching. We see that FacTool’s implementation offers strong accuracy performance on the FEVER datasets, however the full repository required several hundred lines of code, highlighting the development burden of building these applications without abstractions for semantic-bulk processing. By contrast, each LOTUS program offers comparable or higher accuracy, in relatively few lines of code. We also note that FacTool implements its method without batching, whereas LOTUS, by default, leverages batched LM execution for efficiency. To provide an apples-to-apples comparison, we compare FacTool’s un-batched implementation to the LOTUS programs both with and without batching.
@@
First, we see from the Table LABEL:tab:factchecking that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with 6.5 points higher accuracy and lower execution time without batching. The LOTUS-FacTool implementation with batching further decreases execution time compared to the original FacTool implementation by . We see that the next LOTUS pipeline simply changes a single operation, switching the sem_map to a sem_filter, and maintains similar result quality to LOTUS-FacTool while further reducing execution time by in the batched implementation.
@@
Leveraging LOTUS’ filter operation allows the programmer to further optimize the program using model cascades, which increases accuracy by 3 points and increases the batched execution time by compared to LOTUS-FacTool. Figure 27 highlights the diverse performance trade-offs presented the model cascade optimizations used in LOTUS’ filter sem_filter. The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, Llama 70B, or TinyLlama (Zhang et al., 2024), shown by the circles. We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, or TinyLlama and Llama70B to implement the filter casacade. We generate multiple cascade points, shown by the stars, by varying the confidence threshold used. The plot shows that this filter optimization can substantially reduce execution time, and offer diverse accuracy trade-offs, compared to implementing the pipeline with the oracle model, Llama 70B, alone.
@@
Returning to Table LABEL:tab:factchecking, we find that the last LOTUS pipeline (search-map-join), reduces accuracy and increases execution time. Notably the performance trade-offs of different LOTUS programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines. Each LOTUS program we described can be implemented easily in relatively few lines of code, and the best performing one offers higher accuracy and lower execution time than FacTool’s original, un-batched implementation.
@@
| Semantic Similarity Join | 0.106 | 0.120 | 2.91 | 0.00 |
| LOTUS Semantic Join (map-search-filter pattern) | 0.241 | 0.258 | 2,762 | 7,750 |
| LOTUS Semantic Join (nested-loop pattern) | N/A | N/A | 2,144,560* | 6,092,500 |
| LOTUS Semantic Join (search-filter pattern) | 0.155 | 0.186 | 2,640 | 7,500 |
@@
4.2. Application: Extreme Multi-label Classification
4.2.1. Dataset
We evaluate on the Biodex Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and expert-created drug safety reports constructed from each article. The task is to correctly label the drug reactions experience by the patient in each medical article. We sample 250 patient articles for our evaluation. Notably, there are approximately 24,000 possible drug-reaction labels, making this task an extreme multi-label classification task. Due to the large number of possible labels, leveraging an LM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024). We show below that this task can be efficiently modeled and programmed using LOTUS’ semantic join.
4.2.2. Baselines
We consider a simple retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels. We show the pseudocode for this program in Figure 28.
4.2.3. LOTUS Programs
@@
The proposed LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 32. We perform the semantic join using at most 7 map demonstrations and set an LM call budget of 10,000. We use Llama-70b as the LM.
4.2.4. Results
@@
Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while maintaining query efficiency. The table reports the rank-precision@5 (RP@5), rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls required for each program. Since the nested-loop join pattern is prohibitively expensive to run, we show an estimate of the execution time assuming linear scaling in the number of batched calls.
@@
First we compare performance of the LOTUS join program implemented with the map-search-filter pattern to the baseline method, the semantic similarity join. As expected, the LOTUS program offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the retrieval-based similarity join. This highlights the effectiveness of leveraging LMs over the data for complex reasoning-based tasks. We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4 turbo. D’Oosterlinck et al. report 24.73 RP@5 and 27.67 RP@10 for the compiled program, representing comparable result quality to LOTUS’ program, although notably the LOTUS program was not compiled with a prompt optimization system.
@@
We now consider several interesting performance trade-offs presented in the semantic join algorithm. We compare LOTUS map-search-filter join pattern to the naive nested-loop join algorithm and the search-filter join pattern shown in Table 3. We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is untenable and prohibitively costly. By contrast, the search-filter and map-search-filter pattern substantially reduce the LM call budget of the naive algorithm by , using an approximation. While these two approximation patterns have similar efficiency, according to execution time and the number of LM calls, interestingly, they offer substantially different result quality. Specifically, the map-search-filter pattern offers higher RP@5 and higher RP@10 compared to the retrieve-filter pattern on this task. These results highlight the unique opportunities bulk-semantic processing pipelines present for designing new algorithms and optimizations.

### v2

@@ 4.5.1. Gold Algorithm
Our implementation builds on the LM-based summarization pattern studied by prior research works (Wu et al., 2021; Chang et al., 2024; Adams et al., 2023) and deployed systems (lla, [n. d.]; lan, [n. d.]). These implementations primarily leverage one of two aggregation patterns: either a fold pattern, which produces a sequential, linear pass over the data while iteratively updating an accumulated partial answer, or a hierarchical reduce pattern, which recursively aggregates the input data to produce partial answers until a single answer remains. Our aggregation implements the hierarchical pattern, which allows for greater parallelism during query processing and has been shown to produce higher quality results for tasks like summarization in prior work (Chang et al., 2024).
@@ 5. Evaluation
We now evaluate the expressiveness of the semantic operator model as well as the efficiency of our algorithms and optimizations through four diverse applications: fact-checking (Section 5.1), extreme multi-label classification (Section 5.2), search and ranking (Section 5.3), and paper analysis (Section 5.4). We find that the semantic operator model is expressive and easy to program with, capturing state-of-the-art AI pipelines in a few operator calls, and making it easy to express new pipelines that achieve up to higher quality. We also find that our optimizations that take advantage of the declarative nature of semantic operators can accelerate semantic filtering, top-k, group-by and join operators by up to while offering statistical accuracy guarantees. Overall, LOTUS queries match or exceed the accuracy of state-of-the-art AI pipelines for each task while executing up to 28 faster end-to-end.
LOTUS’ current implementation extends the Pandas API (pan, [n. d.]). We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) to support efficient vector search for LOTUS’ semantic search, similarity join and cluster operations.
Unless otherwise stated, we run our local model experiments with 4 A100 GPUs using Llama-3 models (int, [n. d.]), with a batch size of 64 running on vLLM (Kwon et al., 2023). For our experiments that use OpenAI’s GPT-4o-2024-05-13 model (ope, [n. d.]), we control for cost by limiting execution to 64-way thread parallelism. We set temperature to , unless otherwise stated.
@@ 5.1. Application: Fact-Checking
Application. We evaluate on FEVER (Thorne et al., 2018), a claim verification dataset. We use the development dataset, which contains about 38,000 total claims, of which we sample 1,000 for our evaluation. Each claim is labeled with one of three labels, ”Supported”, ”Refuted”, or ”NotEnoughInfo”, and the task is to correctly determine the label of each claim, leveraging evidence from a corpus of million Wikipedia articles. We merge the latter two labels in to a single class, ”Not Supported”, following prior work (Chern et al., 2023) for our evaluation.
Baselines. FacTool (Chern et al., 2023) is a recent research work that proposes a multi-step pipeline for fact-checking involving, claim extraction, query generation, tool querying, evidence collection, and verification. We use FactTool’s open source codebase (gai, [n. d.]) to measure its performance. FactTool’s pipeline, by default, performs retrieval with a Google Search API (cus, [n. d.]). We evaluate the pipeline with both the default retrieval API, and with the open-source ColBERT (Khattab and Zaharia, 2020) index. We find that the results are similar, and we report the results using ColBERT for retrieval to hold the retriever model constant with the implemented LOTUS programs.
LOTUS Programs. We compose several intuitive LOTUS programs, each in less than 50 lines of code. For each one, we use ColBERT as the retriever model for creating the semantic index, and we use Llama-70B as the primary LM, and Llama-8B as the cascade proxy.
@@
| Method | Accuracy | Execution Time (s), batched | Execution Time (s), no batching | LoC |
| FacTool | 80.9 | N/A | 5396.11 | ¿ 750 |
| LOTUS-Factool | 89.9 | 688.90 | 4,454.24 | ¡ 50 |
| LOTUS-fact-filter | 91.2 | 329.1 | 988.95 | ¡ 50 |
| LOTUS-fact-filter (opt.)* | 91.0 | 189.88 | 776.37 | ¡ 50 |
| LOTUS-fact-join | 84.5 | 11,951.35 | 60,364.39 | ¡ 50 |
The next program, LOTUS-fact-filter, makes a simple, single-operator modification to the LOTUS-FacTool program, replacing the semantic map at the end of the program, with a semantic filter. We use 3 demonstrations for the filter. Figure 7 shows the pseudocode for this program. We evaluate this program with and without model cascades applied to the semantic filter.
@@
Results. Table 2 demonstrates the powerful abstraction that semantic operators provide, allowing programmers to quickly write and test programs that compose a few operators to obtain state-of-the-art results. We report the accuracy of each benchmarked method, an estimate of lines of code (LoC), and execution time in seconds both with and without batching averaged over 10 runs. We see that FacTool’s implementation offers strong accuracy performance on the FEVER dataset, however the full repository was written from scratch in several hundred lines of code, representing a significant development burden. By contrast, each LOTUS program offers comparable or higher accuracy, in relatively few lines of code. We also note that FacTool implements its method without batching, whereas LOTUS, by default, leverages batched LM execution for efficiency. To provide an apples-to-apples comparison, we compare FacTool’s un-batched implementation to the LOTUS programs both with and without batching.
@@
First, we see from the Table 2 that LOTUS-FacTool is able to reproduce the result quality and efficiency of the original method’s implementation, with points higher accuracy and lower execution time without batching. The LOTUS-FacTool implementation with batching further decreases execution time compared to FacTool by . We see that the next LOTUS pipeline simply changes a single operation, switching the sem_map to a sem_filter, and maintains similar result quality to LOTUS-FacTool while further reducing execution time by in the batched implementation.
@@
Leveraging LOTUS’ filter operation allows the programmer to further declaratively optimize the program using an approximation, which maintains high accuracy and reduces the batched execution time by compared to LOTUS-FacTool. Figure 9 highlights the diverse performance trade-offs presented by the model cascade optimizations used in LOTUS’ filter sem_filter. The plot shows the accuracy and execution time of the LOTUS-fact-filter pipeline using a single model, either Llama 8b, or Llama 70B,
shown by the circles. We compare this to the performance attainable using a pair of models, Llama 8B and Llama 70B, to implement the filter with our casacade-based approximation. We generate multiple cascade points, shown by the stars, by varying the recall and precision targets of the filter operator, and setting the error probability . The plot shows that this filter optimization can substantially reduce execution time, and offer diverse accuracy trade-offs, compared to implementing the pipeline with the oracle model, Llama 70B, alone. We also verify that the algorithm’s probabilistic guarantees hold by testing 100 trials for recall and precision targets at . We find the targets are met in 99 of the 100 trials, satisfying and exceeding the guaranteed 0.8 success rate.
@@
Returning to Table 2, we find that the last LOTUS pipeline (LOTUS-fact-join), reduces accuracy and increases execution time. Notably the performance trade-offs of different AI programs are non-obvious, highlighting the need for programmable, declarative abstractions so programmers can quickly explore and iterate on their query pipelines. Each LOTUS program we described can be implemented easily in relatively few lines of code, and the best performing one offers higher accuracy and lower execution time than FacTool’s original, un-batched implementation.
5.2. Application: Extreme Multi-label Classification
Dataset. We evaluate on the Biodex Dataset (D’Oosterlinck et al., 2023), which consists of a corpus of biomedical articles, and expert-created drug safety reports constructed from each article. The task is to correctly label the drug reactions experience by the patient in each medical article. We sample 250 patient articles for our evaluation. Notably, there are possible drug-reaction labels, making this task an extreme multi-label classification task. Due to the large number of possible labels, leveraging an LM to perform inference is difficult, and this setting has been studied in prior works (D’Oosterlinck et al., 2024). We show below that this task can be efficiently modeled and programmed using the semantic join.
Baselines We consider a retrieval baseline which uses an E5Model (Wang et al., 2024) as the retriever and performs a semantic-similarity join between the patient articles and the reaction labels.
@@ 5.2. Application: Extreme Multi-label Classification
| Sem-sim-join | 0.106 | 0.120 | 2.91 | 0.00 |
| LOTUS Sem-join** | 0.244 | 0.261 | 5,891.6 | 15,107 |
| nested-loop pattern | N/A | N/A | 2,144,560* | 6,092,500 |
| map-sim-filter pattern** | 0.244 | 0.261 | 5,891.6 | 15,107 |
| sim-filter pattern** | 0.154 | 0.191 | 24,206.8 | 63,724 |
@@
LOTUS Programs. Our LOTUS program performs a semantic join over the drug reaction labels and the medical articles, as shown in Figure 10. We perform the semantic join using at most 7 map demonstrations. We use Llama-70b as the LM and the E5Model (Wang et al., 2024) as the retriever for the semantic index.
@@
Results. Table 3 shows that the proposed LOTUS program makes meaningful traction, obtaining high quality results on this task, while providing query efficiency compared to naive join implementations. The table reports the rank-precision@5 (RP@5) and rank-precision@10 (RP@10), following prior work (D’Oosterlinck et al., 2024), as well as execution time in seconds and the number of LM calls used. Since the nested-loop join pattern is prohibitively expensive to run, we show an estimate of the execution time assuming linear scaling in the number of batched calls.
@@
First we compare performance of the LOTUS join program, declaratively optimized, to the baseline, a semantic similarity join. As expected, the optimized LOTUS join offers substantially higher quality results, with and higher RP@5 and RP@10 respectively compared to the similarity join. This highlights the effectiveness of leveraging LMs over the data for complex reasoning-based tasks. We informally compare these accuracy results to recent work (D’Oosterlinck et al., 2024) which composes a multi-step DSPy (Khattab et al., 2023) program compiled using Llama-2-7b-chat as the student LM and GPT-3.5-turbo as the teacher LM to perform a semantic mapping, followed by a re-ranking step using GPT-4-turbo. D’Oosterlinck et al. report 24.73 RP@5 and 27.67 RP@10 for the compiled program, representing comparable result quality to LOTUS’ program, although notably the LOTUS program was not compiled with a prompt optimization system.
@@
We now consider several interesting performance trade-offs presented by semantic join algorithms. We compare the map-sim-filter join pattern, which was automatically selected in the LOTUS sem-join, to the naive nested-loop join algorithm and the sim-filter join pattern shown in Table 3. We see that nested-loop join pattern, which involves a quadratic LM budget of over 6 million LM calls, is prohibitively costly. By contrast, the sim-filter and map-sim-filter pattern substantially reduce the LM call budget of the naive algorithm by and respectively. LOTUS’ optimizer automatically selects the lower cost algorithm, in this case map-sim-filter. The selected pattern also results in significantly better accuracy due to its higher quality proxy score for this task. Specifically, the map-sim-filter pattern offers higher RP@5 and higher RP@10 compared to the sim-filter.